Traffic signs are regulated by the Manual on Uniform Traffic control Devices MuTcD) The perimeter a rectangular traffic sign is 106 inches Also its length is 15 of inches longer than its width Find the dimensions of this sign.
Answer by Boreal(15235)

Half perimeter is 53 inches, and that equals 1 length and 1 width
w=x
l=x+15
x+x+15=53
2x=38
x=19 in
x+15=34 in
15 x 34 inches